OpenStudy (anonymous):

Multiply. Simplify your answer. Type an exact answer, using radicals as needed.

OpenStudy (anonymous):

\[(4\sqrt{3}-8\sqrt{7})(5\sqrt{3}+10\sqrt{7})\]

OpenStudy (cwrw238):

= 20*3 + 40sqrt3sqrt7 - 40dqrt3sqrt7 - 80*7 = -500

OpenStudy (anonymous):

tyvm